40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Sep 22 01:45:41 2011, Keiko, Update, LSC, Locking status update PRCL-OLT.pdfMICH-OLT.pdf
    Reply  Thu Sep 22 04:49:14 2011, Anamaria, Update, LSC, Locking status update - Some Scripts, No Louck 
    Reply  Thu Sep 22 11:50:37 2011, Koji, Update, LSC, Locking status update 
Message ID: 5512     Entry time: Thu Sep 22 01:45:41 2011     Reply to this: 5513   5516
Author: Keiko 
Type: Update 
Category: LSC 
Subject: Locking status update 

Keiko, Anamaria

Tonight we want to measure the LSC matrix for PRMI and compare the simulation posted last night (#5495).

First. we locked MICH and PRCL, and measured the OLT to see how good the locking is. The following rough swept sine plots are the OLTs for MICH and PRCL. The gain setting was -10 and 0.5 for MICH and PRCL, respectively. Integrators were off. Looking at the measured plots, MICH has about 300 Hz UGF, when the gain is -20, and PRCL has about 300 HZ UGF, too, when the gain is 0.8.

MICH-OLT.pdf

PRCL-OLT.pdf

As these lokings seemed good, so we tried the LSC matrix code written by Anamaria. However it is not working well at this point. When the script add excitations to the exc channels, they kick the optics too much and the lockings are too much disturbed...

Also, we have been trying to lock PRC with the SB resonant, it doesn't work. Looking at the simulated REFL11I (PRCL) signal (you can see it in #5495 too), the CR and SB resonances have the opposite signs... But minus gain never works for PRCL. It only excites the mirror rather than locking.

ELOG V3.1.3-